OpenStudy (anonymous):

a m =0.09 tennis ball initially moving at a speed of 10m\s in the opposite direction of a speed of 18m/s what is the magnitude of the change in momentum of the ball

OpenStudy (anonymous):

First the ball is going in negative direction ?

OpenStudy (anonymous):

the sign is not applicable here

OpenStudy (anonymous):

hello anyone there

OpenStudy (anonymous):

momentum p=mv so first calculate the initial momentum, then the final momentum and then subtract the final momentum from the initial momentum. (provide units!)

OpenStudy (anonymous):

what units are required for this, is it Kgm/s

OpenStudy (anonymous):

i meant mass for the tennis ball. i told u how to do the question now do it.

OpenStudy (anonymous):

the calculation is the easy part the units is where i am loosing grades

OpenStudy (anonymous):

i put Kgm/s for my last answer and they told me i was wrong, i don,t know what other units i can put in

OpenStudy (anonymous):

0.72Kgm/s is my answer for this,

OpenStudy (anonymous):

it is supposed to be like this: initial momentum=m*u =0.09kg*10m/s =? kgm/s final momentum=m*v =0.09kg*(-18)m/s =? kgm/s change in momentum = initial momentum-final momentum =? kgm/s

OpenStudy (anonymous):

i know so now if they tell me i'm, wrong there'll be trouble, thanks for your time

OpenStudy (anonymous):

hey your answer is still wrong!!!! do the calculations correctly

OpenStudy (xishem):

\[\Delta p=p_f-p_i \rightarrow \Delta p=(0.09kg)(10\frac{m}{s})-(0.09kg)(-18\frac{m}{s})=?\]

OpenStudy (xishem):

The sign IS applicable in this case, unlike what you said.